# Assessing the Effects of Monetary Shocks on Macroeconomic Stars: A SMUC-IV Framework

### v1

@@ Abstract
This paper proposes a structural multivariate unobserved components model with external instrument (SMUC-IV) to investigate the effects of monetary policy shocks on key U.S. macroeconomic “stars”—namely, the level of potential output, the growth rate of potential output, trend inflation, and the neutral interest rate. A key feature of our approach is the use of an external instrument to identify monetary policy shocks within the multivariate unobserved components modeling framework. We develop an MCMC estimation method to facilitate posterior inference within our proposed SMUC-IV framework. In addition, we propose an marginal likelihood estimator to enable model comparison across alternative specifications. Our empirical analysis shows that contractionary monetary policy shocks have significant negative effects on the macroeconomic stars, highlighting the nonzero long-run effects of transitory monetary policy shocks.
@@ 1 Introduction
Long-run equilibrium levels of key macroeconomic variables—commonly referred to as “stars”—are central to guiding countercyclical policy and assessing long-run performance. In the conventional view, the macroeconomic stars are unaffected by monetary policy shocks, which are thought to have only transitory effects. Recently, a growing body of evidence challenges this view, showing that monetary shocks can have long-run influences on the economy (e.g., Hanson and Stein, , 2015; Diegel and Nautz, , 2021; Jordà et al., , 2024). A less studied empirical question, however, is how monetary policy shocks affect the macroeconomic stars.
@@
The main objective of this paper is to develop a unified econometric framework to address this question. Multivariate unobserved components models have become the standard econometric tool for analyzing macroeconomic stars (e.g., Kuttner, , 1994; Laubach and Williams, , 2003; Chan et al., , 2016; Zaman, , 2025). These models decompose observed macroeconomic series into long-run trends and short-run cycles, identifying the stars with the trend components. The traditional multivariate unobserved components model framework, albeit useful to model the macroeconomic stars, lacks a formal strategy to identify the effects of monetary policy shocks on the stars.
@@
This paper contributes to the existing literature by introducing a structural multivariate unobserved components model to assess the effects of monetary policy shocks on the macroeconomic stars. The main novelty of our approach is the use of an external instrument to identify monetary policy shocks for structural analysis within a multivariate unobserved components model, which we refer to as SMUC-IV. While the use of external instruments to identify macroeconomic shocks in structural vector autoregressions has become increasingly common (Mertens and Ravn, , 2013; Ramey, , 2016; Stock and Watson, , 2018; Jarociński and Karadi, , 2020; Caldara and Herbst, , 2019; Arias et al., , 2021), very few studies have considered their application within multivariate unobserved components models. The SMUC-IV introduced in this paper addresses this gap. Although the SMUC-IV developed in this paper is used to assess the effects of monetary policy shocks on the macroeconomic stars, the methodology readily extends to broader applications in a straightforward manner.
@@
Our proposed SMUC-IV possesses two desirable characteristics. First, it offers a unified framework for the joint estimation of the macroeconomic stars and the assessment of their responses to monetary policy shocks. This ensures internally consistent estimates of the macroeconomic stars, enabling a coherent analysis of the dynamic interactions between the long-run trends and short-run cyclical fluctuations. Second, our SMUC-IV specification allows for correlations among the innovations of all included trend and cycle components. This flexible correlation structure has been demonstrated to be empirically important in many studies on unobserved components models (Morley et al., , 2003; Basistha and Nelson, , 2007; Grant and Chan, 2017a, ; Grant and Chan, 2017b, ; Hwu and Kim, , 2019).
@@
A further contribution of this paper is the development of an MCMC estimation procedure and a marginal likelihood estimator to facilitate estimation and model comparison across various variants of the proposed SMUC-IV. Specifically, our proposed MCMC method builds upon the precision sampling algorithm for Gaussian state space models developed by Chan and Jeliazkov, (2009). A novel feature of our approach is that it exploits the joint Gaussianity between the observed data and the state parameters when deriving the conditional posterior of the latent states. This allows us to circumvent the need to apply Bayes’ rule, which is particularly cumbersome when the state and measurement equations are correlated (Grant and Chan, 2017a, ; Grant and Chan, 2017b, ; Leiva-Leon and Uzeda, , 2023). Furthermore, one benefit of recognizing the joint Gaussianity between the observed data and the state parameters is that it permits direct derivation of the analytical expression for the likelihood function unconditional on the high-dimensional latent states, which is an essential component of our proposed marginal likelihood estimator. Specifically, in this paper we construct a conditional Monte Carlo improved modified harmonic mean estimator to compute the marginal likelihoods for various specifications of the SMUC-IV. The conditional Monte Carlo method was recently proposed by Chan, (2023) for estimating marginal likelihood for large vector autoregressions, and has been shown to significantly improve estimation accuracy.
@@
In our empirical analysis, we jointly estimate four US macroeconomic stars—the level of potential output, the growth rate of potential output, trend inflation, and the neutral interest rate. For the identification of the monetary policy shock, we employ the orthogonal high-frequency surprise series of Bauer and Swanson, (2023) as our external instrument for monetary policy shocks. The series expands the set of events to include Federal Reserve Chair speeches and orthogonalizes the surprises to pre-announcement variables to strengthen relevance and exogeneity.
@@
To assess the validity of our framework, we conduct a Bayesian model comparison exercise. The results show that the proposed SMUC-IV is more strongly supported by the data than competing specifications. The findings support the relevance of the external instrument for monetary policy shocks, confirm correlations between long-run trends and short-run cycles, and underscore the significant influence of monetary policy shocks on the macroeconomic stars.
@@
We document two main findings from our empirical analysis. First, contractionary monetary policy shocks lead to declines in the growth rate of potential output, trend inflation, and the neutral interest rate. The negative effects on the growth rate of potential output align with the main finding of Jordà et al., (2024) that monetary policy can have long-run effects on the real economy and implies an innovation channel, consistent with a growing literature showing that monetary policy can shape the economy’s long-run productive capacity through its effects on innovation and technological change (e.g., Stadler, , 1990; Moran and Queralto, , 2018; Ma and Zimmermann, , 2023; Fornaro and Wolf, , 2023; Meier and Reinelt, , 2024). Likewise, the decline in trend inflation is consistent with findings that contractionary monetary policy shocks lower long-term inflation expectations (e.g., Jarociński and Karadi, , 2020; Diegel and Nautz, , 2021), suggesting a re-anchoring channel when expectations drift from the inflation target. The decline in the neutral interest rate induced by contractionary monetary policy shocks may operate through two channels: the innovation channel and the re-anchoring channel. Contractionary shocks reduce the growth rate of potential output, and since it is a key driver of the neutral interest rate (see, e.g., Laubach and Williams, , 2003), this decline directly lowers the neutral interest rate. In addition, these shocks reduce trend inflation, which raises the real interest rate and encourages firms to reduce spending on R&D and productive investment—reinforcing the innovation channel and further depressing the neutral interest rate. Second, counterfactual analysis based on historical decompositions shows that, absent these contractionary shocks, the growth rate of potential output (along with the level of potential output), trend inflation, and the neutral interest rate would have been notably higher, implying that monetary policy shocks are important drivers of the stars. Taken together, the results imply that monetary policy can help re-anchor long-run inflation expectations at target, but over-tightening risks lowering the the growth rate of potential output and the neutral interest rate. Our two main findings are robust across a range of alternative specifications.
@@ 2 Econometric Framework
In this section, we introduce our SMUC-IV. Section 2.1 presents the specification of the structural multivariate unobserved components model. Section 2.2 then outlines how the external instruments is employed for structural identification.
@@ 2.1 Model Specification
$r_{t}$ $=\pi_{t}^{*}+r_{t}^{*}+c_{r,t},$ (3)
where $g_{t}$ denotes log level of real output, $\pi_{t}$ denotes inflation, and $r_{t}$ denotes the nominal interest rate. In addition, $g_{t}^{*}$, $\pi_{t}^{*}$, and $r_{t}^{*}$ represent the level of potential output, trend inflation, and the neutral interest rate, respectively. In this paper, we use the real GDP as the real output, GDP deflator inflation as the inflation, and the federal funds effective rate as the nominal interest rate (see Section 5.1 for more details). We assume inflation $\pi_{t}$ and nominal interest rate $r_{t}$ share a common trend component, $\pi_{t}^{*}$ as in Del Negro et al., (2017).
For the trend components, we assume the first difference of potential GDP follows a random walk process as in Grant and Chan, 2017b :
$\Delta g_{t}^{*}$ $=\Delta g_{t-1}^{*}+u_{t}^{g^{*}},$ (4)
where $\Delta g_{t}^{*}=g_{t}^{*}-g_{t-1}^{*}$ denotes the growth rate of potential output. For the trend inflation and neutral interest rate, we assume
@@
The initial conditions $g_{-1}^{*},g_{0}^{*},\pi_{0}^{*}$ and $r_{0}^{*}$ are treated as parameters to be estimated. Our trends are defined, consistent with the Beveridge-Nelson decomposition (Beveridge and Nelson, , 1981), as the infinite-horizon forecast of the actual variables of interest, conditional on the information set available in period $t$, which implies a random walk for the trends and stationary, mean-zero cycles.
The specifications of trend components broadly aligns with unobserved components models used to estimate macroeconomic stars, including strands that focus separately on the level of potential output (e.g., Grant and Chan, 2017a, ; Grant and Chan, 2017b, ), trend inflation (e.g., Chan et al., , 2013, 2018; Mertens, , 2016; Stock and Watson, , 2007, 2016; Hwu and Kim, , 2019; Eo et al., , 2023), and the neutral interest rate (e.g., Laubach and Williams, , 2003; Holston et al., , 2017; Del Negro et al., , 2017).
Regarding the cycle components, let $\mathbf{c}_{t}=(c_{g,t},c_{\pi,t},c_{r,t})^{\prime}$ be a vector of cycle components, and we assume that $\mathbf{c}_{t}$ evolves according to the following VAR($p$) process:
@@
where $\bm{\Phi}_{1},\ldots,\bm{\Phi}_{p}$ are $3\times 3$ autoregressive coefficient matrices and $\mathbf{u}_{t}^{c}$ are the reduced-from innovations. [^1] For simplicity, we set the initial conditions $\mathbf{c}_{0}=\ldots=\mathbf{c}_{1-p}=\mathbf{0}$.
@@
$\bm{\tau}_{t}=\bm{\Psi}_{1}\bm{\tau}_{t-1}+\bm{\Psi}_{2}\bm{\tau}_{t-2}+\mathbf{u}_{t}^{\tau},$ (8)
where $\bm{\Psi}_{1}=\text{diag}(2,1,1)$, $\bm{\Psi}_{2}=\text{diag}(-1,0,0)$ and $\mathbf{u}_{t}^{\tau}=(u_{t}^{g^{*}},u_{t}^{\pi^{*}},u_{t}^{r^{*}})^{\prime}$. Stacking equation (8) over equation (7), our model can be represented as
@@
where $\bm{\eta}_{t}=(\bm{\tau}_{t}^{\prime},\mathbf{c}_{t}^{\prime})^{\prime}$ is a vector containing the trend and cycle components. The coefficient matrices are defined as $\mathbf{A}_{1}=\text{diag}(\bm{\Psi}_{1},\bm{\Phi}_{1})$, $\mathbf{A}_{2}=\text{diag}(\bm{\Psi}_{2},\bm{\Phi}_{2})$, $\mathbf{A}_{i}=\text{diag}(\mathbf{0}_{3},\bm{\Phi}_{i})$ for $i=3,\ldots,p$. The residual vector $\mathbf{u}_{t}=(\mathbf{u}_{t}^{\tau^{\prime}},\mathbf{u}_{t}^{c^{\prime}})$ is of dimension $6\times 1$, which will be described shortly. To complete our model specification, we assume that the residual vector $\mathbf{u}_{t}$ are related to the structural shocks by
@@
where $\bm{\epsilon}_{t}$ is a vector of structural shocks, $\mathbf{B}$ is the contemporaneous response matrix that is assumed to be non-singular, and $\bm{\Sigma}=\mathbf{B}\mathbf{B}^{\prime}$ is the covariance matrix of the residual vector $\mathbf{u}_{t}$. It is known that the contemporaneous response matrix $\mathbf{B}$, hence, the structural shocks $\bm{\epsilon}_{t}$, cannot be separately identified without additional information. In this paper, we employ external instrument to identify monetary policy shock. The next section provides the details of our identification approach.
Our proposed model specified in equations (9)–(10) provides a unified framework for studying the macroeconomic stars and their corresponding cycle components. Furthermore, we allows the innovations of all trend and cycle components, i.e., $\mathbf{u}_{t}$, to be correlated. This contrasts with conventional studies on unobserved components models (e.g., Watson, , 1986; Stock and Watson, , 2007; Chan et al., , 2018; Laubach and Williams, , 2003; Zaman, , 2025), which typically assume these innovations to be independent. Specifically, the contemporaneous response matrix $\mathbf{B}$ is assumed to be non-singular and unrestricted, which implies that $\bm{\Sigma}=\mathbf{B}\mathbf{B}^{\prime}$ is a full covariance matrix. In our empirical analysis, we assess the validity of this full correlation structure through a formal Bayesian model comparison exercise, which indicates strong evidence in favor of this modeling feature supported by the data.

### v2

@@ Abstract
This paper proposes a structural multivariate unobserved components model with external instrument (SMUC-IV) to investigate the effects of monetary policy shocks on key U.S. macroeconomic “stars”—namely, the potential GDP growth, trend inflation, and the neutral interest rate. A key feature of our approach is the use of an external instrument to identify monetary policy shocks within the multivariate unobserved components modelling framework. We develop an MCMC estimation method to facilitate posterior inference within our proposed SMUC-IV framework. In addition, we propose a marginal likelihood estimator to enable model comparison across alternative specifications. Our empirical analysis shows that contractionary monetary policy shocks have negative effects on the macroeconomic stars, highlighting the non-zero long-run effects of transitory monetary policy shocks.
@@ 1 Introduction
Long-run equilibrium levels of key macroeconomic variables—commonly referred to as “stars”—are central for guiding countercyclical policy and assessing long-run performance. In the conventional view, the macroeconomic stars are unaffected by monetary policy shocks. These shocks are typically believed to have only transitory effects. However, recent evidence challenges this view. For instance, jorda2024long show that contractionary monetary policy shocks have negative and long-lasting effects on output. A less studied empirical question, however, is how monetary policy shocks affect the macroeconomic stars. This question is of interest because it provides guidance on how policymakers can bring inflation back to target while supporting sustainable economic growth in the wake of the heightened inflation following the COVID-19 pandemic.
@@
The main objective of this paper is to develop a unified econometric framework to address this question. Multivariate unobserved components models have become the standard econometric tool for analyzing macroeconomic stars (e.g., kuttner1994estimating; laubach2003measuring; chan2016bounded; zaman2025unified). These models decompose observed macroeconomic series into long-run trends and short-run cycles, identifying the stars with the trend components. The traditional multivariate unobserved components model framework, albeit useful to model the macroeconomic stars, lacks a formal strategy to identify the effects of monetary policy shocks on the stars.
@@
This paper contributes to the existing literature by introducing a structural multivariate unobserved components model to assess the effects of monetary policy shocks on the macroeconomic stars. The main novelty of our approach is the use of an external instrument to identify monetary policy shocks for structural analysis within a multivariate unobserved components model, which we refer to as SMUC-IV. While the use of external instruments to identify macroeconomic shocks in structural vector autoregressions has become increasingly common (mertens2013dynamic; ramey2016macroeconomic; stock2018identification; jarocinski2020deconstructing; caldara2019monetary; arias2021inference), very few studies have considered their application within multivariate unobserved components models. The SMUC-IV introduced in this paper addresses this gap. Although the SMUC-IV developed in this paper is used to assess the effects of monetary policy shocks on the macroeconomic stars, the methodology readily extends to broader applications in a straightforward manner.
@@
Our proposed SMUC-IV possesses two desirable characteristics. First, it offers a unified framework that jointly estimates the macroeconomic stars and assesses their responses to monetary policy shocks, rather than focusing on a single star. This ensures internally consistent estimates of the macroeconomic stars, enabling a coherent analysis of the dynamic interactions between the long-run trends and short-run cyclical fluctuations. Second, our SMUC-IV specification allows for correlations among the innovations of all included trend and cycle components. This flexible correlation structure has been shown to be empirically important in many studies on unobserved components models (morley2003beveridge; basistha2007new; grant2017bayesian; grant2017reconciling; hwu2019estimating).
@@
A further contribution of this paper is the development of an MCMC estimation procedure and a marginal likelihood estimator, designed to facilitate estimation and model comparison across various variants of the proposed SMUC-IV. Specifically, our proposed MCMC method builds upon the precision sampling algorithm for Gaussian state space models developed by chan2009efficient. A novel feature of our approach is that it exploits the joint Gaussianity between the observed data and the state parameters, which allows for the direct derivation of the conditional posterior of the states. This approach bypasses the need for separate computations of the conditional likelihood and prior, leading to a more straightforward implementation, particularly when the state and measurement equations are correlated (grant2017bayesian; grant2017reconciling; leiva2023endogenous). A related benefit of recognizing the joint Gaussianity between the observed data and the state parameters is that it permits direct derivation of the likelihood function unconditional on the high-dimensional latent states. This unconditional likelihood is an essential component for our proposed marginal likelihood estimator. Specifically, we construct a conditional Monte Carlo improved modified harmonic mean estimator to compute the marginal likelihoods for various specifications of the SMUC-IV. The conditional Monte Carlo method was recently proposed by chan2023comparing for estimating marginal likelihood for large vector autoregressions, and has been shown to significantly improve estimation accuracy.
@@
In our empirical analysis, we jointly estimate three US macroeconomic stars—the potential GDP growth, trend inflation, and the neutral interest rate. For the identification of the monetary policy shock, we employ the orthogonal high-frequency surprise series of bauer2023reassessment as our external instrument for monetary policy shocks. The series expands the set of events to include Federal Reserve Chair speeches and orthogonalizes the surprises to pre-announcement variables to strengthen relevance and exogeneity.
@@
To assess the validity of our framework, we conduct a Bayesian model comparison exercise in our empirical analysis. The results show that the proposed SMUC-IV is more strongly supported by the data than competing specifications. The findings support the relevance of the external instrument for monetary policy shocks, confirm correlations between long-run trends and short-run cycles, and—most importantly—underscore the influence of monetary policy shocks on the macroeconomic stars.
@@
We document two main findings on how monetary policy shocks affect the macroeconomic stars. First, contractionary monetary policy shocks lead to declines in the macroeconomic stars. The negative effects on potential GDP growth are consistent with the empirical evidence in moran2018innovation, garga2021output, ma2023monetary, jorda2024long, and meier2024monetary, which shows that tighter monetary policy dampens investment in innovation and, in turn, productivity. Likewise, the decline in the neutral interest rate following contractionary monetary policy shocks may be interpreted as reflecting the same adverse effects of monetary tightening on investment in innovation. The fall in trend inflation is consistent with findings that contractionary monetary policy shocks lower long-term inflation expectations (e.g., jarocinski2020deconstructing; diegel2021long). Second, counterfactual analysis based on historical decompositions shows that, in the absence of these contractionary shocks, potential GDP growth, trend inflation, and the neutral interest rate would have been notably higher, implying that monetary policy is a driving force in shaping the stars. Our two main findings are robust across a range of alternative specifications.
The negative effects of contractionary monetary policy shocks on potential GDP growth and the neutral interest rate can be rationalized by the mechanism highlighted in the Keynesian endogenous growth model with nominal rigidities in fornaro2023scars, which shows that monetary tightening may trigger a “supply–demand doom loop”: lower demand reduces firms’ profits, which curtails investment; weaker investment slows expected productivity growth, lowers household wealth, and further depresses demand—ultimately leading to substantial declines in potential GDP growth and the neutral interest rate. The fall in trend inflation may reflect a re-anchoring channel, whereby expectations that have drifted away from the inflation target are deliberately pulled back, as discussed in detail in diegel2021long. The results imply that central banks face a dilemma: disinflating the economy may come at the cost of lower potential GDP growth and a lower neutral interest rate. This, in turn, points to a potential role for fiscal interventions that support business investment and the economy’s productive capacity during periods of disinflation.
@@ 2 Econometric Framework
In this section, we introduce our SMUC-IV. Section 2.1 presents the specification of the structural multivariate unobserved components model. Section 2.2 then outlines how the external instrument is employed for structural identification.
@@ 2.1 Model Specification
$i_{t}$ $=\pi_{t}^{*}+r_{t}^{*}+c_{i,t},$ (3)
where $g_{t}$ denotes real GDP growth, $\pi_{t}$ denotes inflation, and $i_{t}$ denotes the nominal interest rate. In addition, $g_{t}^{*}$, $\pi_{t}^{*}$, and $r_{t}^{*}$ represent the potential GDP growth, trend inflation, and the neutral interest rate, respectively. In this paper, we use the log difference of real GDP as the real GDP growth, GDP deflator inflation as the inflation, and the federal funds effective rate as the nominal interest rate (see Section 5.1 for more details). We assume inflation $\pi_{t}$ and nominal interest rate $i_{t}$ share a common trend component, $\pi_{t}^{*}$ as in del2017safety.
We assume the trend components of real GDP growth, inflation and real interest rate to follow random walk processes:
$g_{t}^{*}$ $=g_{t-1}^{*}+u_{t}^{g^{*}},$ (4)
@@
The initial conditions $g_{0}^{*},\pi_{0}^{*}$ and $r_{0}^{*}$ are treated as parameters to be estimated. Our trends are defined, consistent with the Beveridge-Nelson decomposition (beveridge1981new), as the infinite-horizon forecast of the actual variables of interest, conditional on the information set available in period $t$, which implies a random walk for the trends and stationary, mean-zero cycles.
The specifications of trend components broadly aligns with unobserved components models used to estimate macroeconomic stars, including strands that focus separately on the potential GDP growth (e.g., grant2017bayesian; grant2017reconciling), trend inflation (e.g., chan2013new; chan2018new; mertens2016measuring; stockwatson2; stock2016core; hwu2019estimating; eo2023understanding), and the neutral interest rate (e.g., laubach2003measuring; holston2017measuring; del2017safety).
Regarding the cycle components, let $\mathbf{c}_{t}=(c_{g,t},c_{\pi,t},c_{i,t})^{\prime}$ be a vector of cycle components, and we assume that $\mathbf{c}_{t}$ evolves according to the following VAR($p$) process:
@@
where $\bm{\Phi}_{1},\ldots,\bm{\Phi}_{p}$ are $3\times 3$ autoregressive coefficient matrices and $\mathbf{u}_{t}^{c}$ are the reduced-form innovations. [^1] For simplicity, we set the initial conditions $\mathbf{c}_{0}=\ldots=\mathbf{c}_{1-p}=\mathbf{0}$.
@@
$\bm{\tau}_{t}=\bm{\tau}_{t-1}+\mathbf{u}_{t}^{\tau},$ (8)
where $\mathbf{u}_{t}^{\tau}=(u_{t}^{g^{*}},u_{t}^{\pi^{*}},u_{t}^{r^{*}})^{\prime}$. Stacking equation (8) over equation (7), our model can be represented as
@@
where $\bm{\eta}_{t}=(\bm{\tau}_{t}^{\prime},\mathbf{c}_{t}^{\prime})^{\prime}$ is a vector containing the trend and cycle components. The coefficient matrices are $\mathbf{A}_{1}=\text{diag}(\mathbf{I}_{3},\bm{\Phi}_{1})$, $\mathbf{A}_{i}=\text{diag}(\mathbf{0}_{3\times 3},\bm{\Phi}_{i})$ for $i=2,\ldots,p$. The residual vector $\mathbf{u}_{t}=(\mathbf{u}_{t}^{\tau^{\prime}},\mathbf{u}_{t}^{c^{\prime}})$ is of dimension $6\times 1$, which will be described shortly. To complete our model specification, we assume that the residual vector $\mathbf{u}_{t}$ is related to the structural shocks by
@@
where $\bm{\epsilon}_{t}$ is a vector of structural shocks, $\mathbf{B}$ is the contemporaneous response matrix that is assumed to be non-singular, and $\bm{\Sigma}=\mathbf{B}\mathbf{B}^{\prime}$ is the covariance matrix of the residual vector $\mathbf{u}_{t}$. It is known that the contemporaneous response matrix $\mathbf{B}$, hence, the structural shocks $\bm{\epsilon}_{t}$, cannot be separately identified without additional information. In this paper, we employ an external instrument to identify monetary policy shocks. The next section provides the details of our identification approach.
Our proposed model specified in equations (9)–(10) provides a unified framework for studying the macroeconomic stars and their corresponding cycle components. Furthermore, we allow the innovations of all trend and cycle components, i.e., $\mathbf{u}_{t}$, to be correlated. This contrasts with conventional studies on unobserved components models (e.g., watson1986univariate; stockwatson2; chan2018new; laubach2003measuring; zaman2025unified), which typically assume these innovations to be independent. Specifically, the contemporaneous response matrix $\mathbf{B}$ is assumed to be non-singular and unrestricted, which implies that $\bm{\Sigma}=\mathbf{B}\mathbf{B}^{\prime}$ is a full covariance matrix. In our empirical analysis, we assess the validity of this full correlation structure through a formal Bayesian model comparison exercise, which indicates strong evidence in favor of this modeling feature supported by the data.